OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x − y = 6 4x − 3y = 12

OpenStudy (anonymous):

multiplica la primera ecuación por -3 a ambos miembros

OpenStudy (anonymous):

-3x +3y=18 4x - 3y=12

jimthompson5910 (jim_thompson5910):

you want to take advantage of the fact that x - x = 0 for any x value So you want to apply this to the y variable Notice how adding -y to -3y gives you -4y, which is not 0. So we need to turn -y into +3y, and we do this by multiplying the top equation by -3 (both sides) to get -3x +3y=18 Now notice how 3y+(-3y) = 3y-3y = 0 So the first step would be to multiply both sides of the first equation by -3.